# Hypergrah-Enhanced Dual Convolutional Network for Bundle Recommendation

### v1

@@ IV-B1 Embedding Initialization Layer
IV-B2 Dual Convolutional Layer
In a graph, messages are passed from nodes to edges and then from edges back to nodes[17][18]. Message passing is a paradigm for graph neural networks, specifically the propagation of features from one node to neighboring nodes. Similarly, in hypergraphs, messages are typically passed from nodes to hyperedges and then back to nodes. Although Feng et al. [19] proposed a hypergraph convolutional neural network, this network structure is not specific to recommender systems and does not take into account the characteristics of recommender systems. He et al. [20] demonstrated that, in recommender systems, feature transformations and nonlinear activations affect the model’s performance. Based on the previous work, the Efficient hypergraph convolution was designed in work [14].
@@ IV-B2 Dual Convolutional Layer
Although this design simplifies the hypergraph convolution operation and lowers the training cost, it neglects the features of bundle recommendation. The complete hypergraph introduces item elements, item-user interactions, and item-bundle interactions. Meanwhile, graph convolution operations on the complete hypergraph exploit auxiliary information but do not possess the module to return the auxiliary information for the task of bundle recommendation. We included $E_{i}$ to improve the learning of $E_{u}$ and $E_{b}$, yet the personalized scoring Equation 8 for the model does not feature $E_{i}$. As a result, we require new modules to enhance the expression of both $E_{u}$ and $E_{b}$, so that the features regarding an item in $E_{u}$ and $E_{b}$ will ultimately predict user-bundle interactions.
Therefore, we design the following convolution operation: the $l$-th layer embedding vector $E^{(l)}$ is computed as ${\bar{E}}^{(l)}$ after graph convolution on the complete hypergraph. We extract the embedding vectors of users in ${\bar{E}}^{(l)}$, ${\bar{E}}^{(l)}_{u}$ and bundles embedding vectors ${\bar{E}}^{(l)}_{b}$. ${\bar{E}}^{(l)}_{u}$ and ${\bar{E}}^{(l)}_{b}$ are respectively performed again on U-B graphs $h_{ub}$. The graph convolution is computed to obtain ${\tilde{E}}^{(l)}_{u}$ and ${\tilde{E}}^{(l)}_{b}$, while the output embedding $E^{(l+1)}$ is obtained by summing ${\bar{E}}^{(l)}$ with ${\tilde{E}}^{(l)}_{u}$ and ${\tilde{E}}^{(l)}_{b}$ according to parameters $\alpha$ and $\beta$. Since the item’s embedding does not participate in the convolution computation on the U-B graph, ${\bar{E}}^{(l)}_{i}$ is equal to $E_{i}^{(l+1)}$. Assuming that the complete hypergraph constructed is $H$, the graph consisting of the purchase interactions $h_{ub}$ is $h$. Then the convolution operation on the bipartite graph that we designed can be expressed by Equation 6.
@@
Since the model we designed contains multiple layers, only considering the embedding of the last layer will neglect the message of the neighbors mined by other layers. However, directly connecting the embedding of different layers will increase the model’s complexity. To address this problem, we draw inspiration from the method proposed in LightGCN [20], which combines the embedding of different layers by decaying them. This approach strategically integrates the embeddings from various layers, enabling the preservation and utilization of information mined at each level. Crucially, it achieves this without increasing the dimensionality of the embedding vectors, thereby maintaining the computational efficiency. And the combination method can be expressed by the formula, which is the final representation of users and bundles.
@@ IV-B4 Model Training
where $E_{u}$in $R^{d}$ denotes the embedding vector of user $u$ and $W$in $R_{d}$ is the learnable parameter. The UIB loss is a mixture of point-by-point loss and pairwise loss, which can be formulated as follows:
@@
V EXPERIMENTS
In this section, we delineate the comprehensive evaluation setup for our proposed approach, which we have named ‘HED’. The evaluation is meticulously designed to rigorously assess the efficacy, efficiency, and overall performance of ‘HED’ in various scenarios and against multiple benchmarks.
V-A Experiments Setup
Datasets. We choose the two most common datasets in bundle recommendation from different domains in the real world, Netease [^5] https://music.163.com/ and Youshu [^6] https://www.yousuu.com/, to validate our proposed model. Youshu is a book website from China, and the dataset provides interaction data of users, books, and bundles containing a series of books. The Netease dataset is constructed by crawling Netease Cloud Music, and it provides interaction data of users, music, and bundles containing a series of music. The parameters of the two datasets are shown in Table I.
TABLE I: Dataset statistics
@@ V-A Experiments Setup
Baselines. In the subsequent phase of our experiments, we conduct a comprehensive comparison between ‘HED’ and a selection of representative baseline methods. This comparative analysis is essential to validate the effectiveness and superior performance of ‘HED’. To ensure a fair and thorough comparison, we have chosen baseline methods that are widely recognized and have demonstrated robust performance in similar contexts. These methods encompass a range of techniques and algorithms, providing a diverse benchmark for assessing the strengths and potential areas of improvement for ‘HED’.
Next, we present these baseline methods:
@@
  • BundleGT[15] is a graph transformer model that contains a token embedding layer, a hierarchical graph transformer layer, and a prediction layer
Training Details. In our experimental framework, we allocated 80% of the collected data for training ‘HED’, while the remaining 20% served as the validation set to assess the model’s performance. All experiments were carried out on a uniform hardware setup, featuring an Intel(R) Xeon(R) Platinum 8336C CPU and an Nvidia GeForce RTX 4090 GPU, ensuring consistency in computational performance across tests. Given the distinct characteristics of our datasets, we tailored the experimental parameters to optimize performance for each. Details are listed in Table II.
TABLE II: The parameter details.
NetEase Youshu
@@
$L_{2}$-norm 0.1
V-B Comparative Results
In our comparative analysis, we endeavored to reproduce the baseline methods using publicly available source code or by closely adhering to the methodologies outlined in their respective publications. This approach was taken to ensure a fair and accurate comparison with our ‘HED’ model. However, it’s important to note that due to hardware constraints, the method proposed in Wei et al. (2023) [15] could not be effectively executed on the Netease dataset. Consequently, this limitation has resulted in the absence of data for Recall@80 and NDCG@80 metrics for this particular method. Table III presents a comprehensive performance comparison between ‘HED’ and the baseline methods.
TABLE III: Performance comparisons on two real-world datasets with the baselines.
@@ V-B Comparative Results
HED 0.3153 0.1903 0.423 0.2197 0.5377 0.2452
%Improv. 3.41% 4.50% 1.54% 3.53% 1.30% 2.90%
HED has become the new state-of-the-art model for bundled recommendation. According to table III, our proposed method is a generalized approach that outperforms the baseline methods of Youshu and NetEase by an average of 2.86% and 5.8%. We attribute this to the complete hypergraph’s superior ability to represent the ternary relationship between users, items, and bundles, and the dual convolution module’s capacity to extract information embedded in this relationship. This module also enhances the representation of users and bundles by leveraging information from the U-B graph, which significantly improves bundle recommendation task performance.
Through the joint analysis of the results of two datasets, we find that the improvements on NetEase are better than those on Youshu. The disparity in dataset size and interaction density between the Youshu and NetEase datasets could account for the observed differences. Specifically, the Youshu dataset has fewer users, items, and bundles than the NetEase dataset, as well as denser interactions among them, as indicated in Table I. This leads to the model’s $Recall@k$ and $NDCG@k$ metrics showing a significant improvement on the Youshu dataset compared to the NetEase dataset, resulting in a relatively smaller increase in performance on the former. However, in real-life datasets, there is often a similarity to the NetEase dataset where the number of users, items, and bundles is vast, and the interaction data between them is sparse. The proposed HED is applicable to practical situations.
The dual convolution module enhances the precision of the package proposition. Specifically, although HED performs better than previous models in both $Recall@k$ and $NDCG@k$, the improvement of $NDCG@k$ for the recommendation accuracy measure is greater than that of $Recall@k$, the maximum difference of 1.99%, the minimum difference of 0.05%. This may be due to the fact that previous work does not revert to the recommendation task after extracting information regarding the item. In contrast, our dual convolution module, which is designed to mine complete hypergraph information, can enhance the representation of users and bundles based on U-B graph information, thus improving the recommendation accuracy through the final personalized bundle recommendation score Formula 8.
V-C Ablation Study
@@
TABLE IV: Ablated models analysis on Youshu dataset.
Recall@20 NDCG@20 Recall@40 NDCG@40 Recall@80 NDCG@80
HED 0.316 (100%) 0.1892 (100%) 0.4209 (100%) 0.2175 (100%) 0.5395 (100%) 0.2445 (100%)
HED-c 0.3086 (97.66%) 0.1841 (97.3%) 0.4164 (98.93%) 0.2128 (97.84%) 0.5304 (98.31%) 0.2397 (98.04%)
HED-cu 0.2987 (94.53%) 0.1776 (93.87%) 0.4044 (96.08%) 0.2057 (94.57%) 0.5235 (97.03%) 0.2323 (95.01%)
HED-cb 0.2965 (93.83%) 0.1779 (94.03%) 0.4025 (95.63%) 0.2064 (94.90%) 0.5267 (97.63%) 0.2337 (95.58%)
HED-cbu 0.3011 (95.28%) 0.1792 (94.71%) 0.4046 (96.13%) 0.2076 (95.45%) 0.5254 (97.39%) 0.2347 (95.99%)
We summarize the results of our experiments in Table IV, where the leading number is the performance of the model, and the trailing percentage is the current performance as a percentage of the performance of the full model. With the table  IV we can get the following conclusions:
@@
V-D Sensitivity Analysis
@@
Figure 4: Recall and NDCG performance under different $d$
Figure 5: Recall and NDCG performance under different $n$.
Figure 6: Recall and NDCG performance under different $\alpha$.
Figure 7: Recall and NDCG performance under different $\beta$
  • $d$ represents the dimension, which is the size of the embedding vectors for users, items, and bundles. The embedding vectors have a size of 64 for the baseline. The findings from the sensitivity analysis regarding the parameter $d$ are outlined in Figure 4, the left-hand line graph shows the recall metrics for different $d$ values, and the right-hand line graph shows the NDCG metrics for different $d$ values. When $d=16$, there is a more significant decline in model performance than the baseline. This is because, when $d$ is too small, the embedding vectors are insufficient in expressing the complete information acquired from the hypergraph and U-B graph. On the other hand, when $d>=32$, the embedding vectors exceed the dimensionality limit and can completely express the information derived from the complete hypergraph and U-B graph. The experiments support the claim that for $d>=32$, the model’s performance is similar to that of the baseline. Furthermore, the model performs better in terms of individual metrics. For instance, with $d=32$, all three metrics - $Recall@40$, $NDCG@40$, and $NDCG@80$ outperformed the baseline metrics.
@@
VI Conclusions
In this study, the significance of creating user-item-bundle triples in bundle recommendation is examined. A comprehensive hypergraph construction approach and HED model are introduced. A comprehensive hypergraph is a technique for conveying the ternary relationship formed by user, item, and bundle interactions. The HED model includes an embedding initialization layer, a dual convolutional layer, and a prediction layer. The embedding initialization layer produces embedding vectors for every user, item, and bundle. The dual convolutional layer extracts complete hypergraph information and enhances the embedding vectors by merging the U-B graph information. The prediction layer provides recommendations based on the final embedding vectors. A substantial volume of experiments conducted on Youshu and NetEase empirically prove the superiority of the HED model we have proposed.

### v2

@@ 4.2.1 Embedding initialization layer
4.2.2 Dual convolutional layer
In a graph, messages are passed from nodes to edges and then from edges back to nodes[kipf2016semi][38]. Message passing is a paradigm for graph neural networks, specifically the propagation of features from one node to neighboring nodes. Similarly, in hypergraphs, messages are typically passed from nodes to hyperedges and then back to nodes. Although Feng et al. [12] proposed a hypergraph convolutional neural network, this network structure is not specific to recommender systems and does not take into account the characteristics of recommender systems. He et al. [17] demonstrated that, in recommender systems, feature transformations and nonlinear activations affect the model’s performance. Based on the previous work, the Efficient hypergraph convolution was designed in work [49].
@@ 4.2.2 Dual convolutional layer
Although this design simplifies the hypergraph convolution operation and reduces the training cost, it ignores the characteristics of bundled recommendation. The complete hypergraph introduces item elements, item-user interactions, and item-bundle interactions. Meanwhile, graph convolution operations on complete hypergraphs use auxiliary information, but do not have modules that focus on bundle recommendations. We include $E_{i}$ to improve the learning of $E_{u}$ and $E_{b}$, but the model’s personalised scoring Equation (7) does not include $E_{i}$. As a result, we need new modules that utilise $E_{i}$ to enhance the expressiveness of $E_{u}$ and $E_{b}$ so that $E_{u}$ and $E_{b}$ can more accurately predict the user’s interactions with the bundle. Therefore, we design the following convolution operation: the $l$-th layer embedding vector $E^{(l)}$ is computed as ${\bar{E}}^{(l)}$ after graph convolution on the complete hypergraph. We extract the embedding of users in ${\bar{E}}^{(l)}$, ${\bar{E}}^{(l)}_{u}$ and bundles embedding ${\bar{E}}^{(l)}_{b}$. ${\bar{E}}^{(l)}_{u}$ and ${\bar{E}}^{(l)}_{b}$ are respectively performed again on U-B graphs $h_{ub}$. The graph convolution is computed to obtain ${\tilde{E}}^{(l)}_{u}$ and ${\tilde{E}}^{(l)}_{b}$, while the output embedding $E^{(l+1)}$ is obtained by summing ${\bar{E}}^{(l)}$ with ${\tilde{E}}^{(l)}_{u}$ and ${\tilde{E}}^{(l)}_{b}$ according to parameters $\alpha$ and $\beta$. Since the item’s embedding does not participate in the convolution computation on the U-B graph, ${\bar{E}}^{(l)}_{i}$ is equal to $E_{i}^{(l+1)}$. Assuming that the complete hypergraph constructed is $H$, the graph consisting of the purchase interactions $h_{ub}$ is $h$. Then the convolution operation on the bipartite graph that we designed can be expressed by Equation (5).
@@
Since the model we designed contains multiple layers, only considering the embedding of the last layer will neglect the message of the neighbors mined by other layers. However, directly connecting the embedding of different layers will increase the model’s complexity. To address this problem, we draw inspiration from the method proposed in LightGCN [17], which combines the embedding of different layers by decaying them. This approach strategically integrates the embeddings from various layers, enabling the preservation and utilization of information mined at each level. Crucially, it achieves this without increasing the dimensionality of the embedding, thereby maintaining the computational efficiency. The combination method can be expressed by the Equation (6), which is the final representation of users and bundles.
@@ 4.2.4 Model Training
where $E_{u}$in $R^{d}$ denotes the embedding vector of user $u$ and $W$in $R^{d}$ is the learnable parameter. The UIB loss is a mixture of point-by-point loss and pairwise loss, which can be formulated as follows:
@@
Table 1: Dataset statistics
@@
5 EXPERIMENTS
In this section, we delineate the comprehensive evaluation setup for HED. The evaluation is meticulously designed to rigorously assess the efficacy, efficiency, and overall performance of HED in various scenarios and against multiple benchmarks.
5.1 Experimental Setup
Datasets. We choose the two most common datasets in bundle recommendation from different domains in the real world, Netease [^5] https://music.163.com/ and Youshu [^6] https://www.yousuu.com/, to validate our proposed model. The Netease dataset is constructed by crawling Netease Cloud Music, and it provides interaction data of users, music, and bundles containing a series of music. Youshu is a book website from China, and the dataset provides interaction data of users, books, and bundles containing a series of books. The parameters of the two datasets are shown in Table 1.
@@ 5.1 Experimental Setup
Baselines. In the subsequent phase of our experiments, we conduct a comprehensive comparison between HED and a selection of representative baseline methods. This comparative analysis is essential to validate the effectiveness and superior performance of HED. To ensure a fair and thorough comparison, we have chosen baseline methods that are widely recognized and have demonstrated robust performance in similar contexts. These methods encompass a range of techniques and algorithms, providing a diverse benchmark for assessing the strengths and potential areas of improvement for HED. In specific, we compare HED with the following baseline methods:
@@
  • BundleGT[42] is a graph transformer model that contains a token embedding layer, a hierarchical graph transformer layer, and a prediction layer.
Training Details. In our experimental framework, we allocated 80% of the collected data for training HED, while the remaining 20% served as the validation set to assess the model’s performance. All experiments were carried out on a uniform hardware setup, featuring an Intel(R) Xeon(R) Platinum 8336C CPU and an Nvidia GeForce RTX 4090 GPU, ensuring consistency in computational performance across tests. Given the distinct characteristics of our datasets, we tailored the experimental parameters to optimize performance for each. A search was conducted to identify an appropriate learning rate within the set range 5e-1, 5e-2, 5e-3, 5e-4, 5e-5, and the rate of 5e-3 was selected for further experimentation. Details are listed in Table 2.
5.2 Comparative Results
In our comparative analysis, we endeavored to reproduce the baseline methods using publicly available source code or by closely adhering to the methodologies outlined in their respective publications. This approach was taken to ensure a fair and accurate comparison with our HED model. However, it’s important to note that due to hardware constraints, the method proposed in Wei et al. (2023) [42] could not be effectively executed on the Netease dataset. Consequently, this limitation has resulted in the absence of data for Recall@80 and NDCG@80 metrics for this particular method.
Table 2: The parameter details.
Parameter NetEase Youshu
@@ 5.2 Comparative Results
L2-norm 0.1
Table 3 presents a comprehensive performance comparison between HED and the baseline methods. In the table, the best results for the comparison methods are underlined, while the best progress for HED is shown in bold. Additionally, the percentage improvement is calculated. The number after HED represents the size of the embeddings. The experimental setup is identical to that described in Table 2, with the exception that HED-32 has been trained for 500 epochs on the Youshu dataset.
Table 3: Performance comparisons on two real-world datasets with the baselines.
@@
HED-32 0.3157 0.1886 0.4223 0.2175 0.5412 0.2445
HED-64 0.3165 0.1904 0.4244 0.2197 0.5427 0.2466
HED-128 0.3122 0.1885 0.421 0.2176 0.5373 0.2438
%Improv. 3.80% 4.56% 1.87% 3.53% 2.24% 3.48%
HED has become the new state-of-the-art model for bundled recommendation. According to Table 3, our proposed method is a generalized approach that outperforms the baseline methods of Youshu and NetEase by an average of 3.25% and 6.26%. We attribute this to the complete hypergraph’s superior ability to represent the ternary relationship between users, items, and bundles, and the dual convolution module’s capacity to extract information embedded in this relationship. This module also enhances the representation of users and bundles by leveraging information from the U-B graph, which significantly improves bundle recommendation task performance.
Through the joint analysis of the results of two datasets, we find that the improvements on NetEase are better than those on Youshu. The disparity in dataset size and interactions density between the Youshu and NetEase datasets could account for the observed differences. Specifically, the Youshu dataset has fewer users, items, and bundles than the NetEase dataset, as well as denser interactions among them, as indicated in Table 1. This leads to the model’s $Recall@k$ and $NDCG@k$ metrics showing a significant improvement on the Youshu dataset compared to the NetEase dataset, resulting in a relatively smaller increase in performance on the former. However, in real-life datasets, there is often a similarity to the NetEase dataset where the number of users, items, and bundles is vast, and the interaction data between them is sparse. The proposed HED is applicable to practical situations.
The dual convolution module enhances the precision of the bundle proposition. Specifically, although HED performs better than previous models in both $Recall@k$ and $NDCG@k$, the improvement of $NDCG@k$ for the recommendation accuracy measure is greater than that of $Recall@k$, the maximum difference of 1.66%, the minimum difference of 0.04%. This may be due to the fact that previous work does not revert to the recommendation task after extracting information regarding the item. In contrast, our dual convolution module aims to mine the complete hypergraph information, which can enhance the representation of users and bundles based on U-B graph information to improve recommendation accuracy. In addition, HED has the highest performance improvement in the Top 20 compared to the Top 40, and the Top 80 also proves the above point.
5.3 Ablation Study
@@
Table 4: Ablated models analysis on Youshu dataset.
HED HED-c HED-cu HED-cb HED-cbu
Recall@20 0.3165 (100%) 0.3086 (96.94%) 0.2987 (94.38%) 0.2965 (93.68%) 0.3011 (95.13%)
NDCG@20 0.1904 (100%) 0.1841 (96.69%) 0.1776 (93.28%) 0.1779 (93.43%) 0.1792 (94.12%)
Recall@40 0.4244 (100%) 0.4164 (98.11%) 0.4044 (95.29%) 0.4025 (94.84%) 0.4046 (95.33%)
NDCG@40 0.2197(100%) 0.2128 (96.86%) 0.2057 (93.63%) 0.2064 (93.95%) 0.2076 (94.49%)
Recall@80 0.5427 (100%) 0.5304 (97.73%) 0.5235 (96.46%) 0.5267 (97.05%) 0.5254 (96.81%)
NDCG@80 0.2466 (100%) 0.2397 (97.20%) 0.2323 (94.20%) 0.2337 (94.77%) 0.2347 (95.17%)
We summarize the results of our experiments in Table 4, where the leading number is the performance of the model, and the trailing percentage is the current performance as a percentage of the performance of the full model. With the table  4 we can get the following conclusions:
Figure 4: Recall and NDCG performance under different $d$
Figure 5: Recall and NDCG performance under different $n$
Figure 6: Recall and NDCG performance under different $\alpha$.
@@
5.4 Sensitivity Analysis
Figure 7: Recall and NDCG performance under different $\beta$
@@ 5.4 Sensitivity Analysis
  • $d$ represents the dimension, which is the size of the embedding for users, items, and bundles. The embedding has a size of 64 for the baseline. The findings from the sensitivity analysis regarding the parameter $d$ are outlined in Figure 4, the left-hand line graph shows the recall metrics for different $d$ values, and the right-hand line graph shows the NDCG metrics for different $d$ values. When $d=16$, there is a more significant decline in model performance than the baseline. This is because, when $d$ is too small, the embedding are insufficient in expressing the complete information acquired from the hypergraph and U-B graph. On the other hand, when $d>=32$, the embedding exceeds the dimensionality limit and can completely express the information derived from the complete hypergraph and U-B graph. The experiments support the claim that for $d>=32$, the model’s performance is similar to that of the baseline. Furthermore, the model performs better in terms of individual metrics. For instance, with $d=32$, all three metrics - $Recall@40$, $NDCG@40$, and $NDCG@80$ outperformed the baseline metrics.
@@
5.5 Discussion on $h_{ii}$ in hypergraphs
In order to better study our proposed complete hypergraph, we conducted experiments under the Youshu dataset for the case where $h_{ii}$ is the unit matrix.
The experimental results are shown in Table 5, where although the performance of $Recall@80$ improves, it performs poorly under the more important Top $20$ and Top $40$ items, and there is a significant dip in recommendation accuracy. This may be due to the fact that the number of items in the datasets is more than a multiple of the number of users and bundles combined ($2.56$ times in the YouShu dataset, $2.99$ times in the NetEase dataset). If $h_{ii}$ is set to a non-zero matrix, it will cause the constructed complete hypergraph to contain too much item information which will affect the quality of the embedding of users and bundles, and affect the final results. This is why it performs better than the original model only in the $Recall@80$ metric.
Table 5: Experiments with $h_{ii}$ on the Youshu dataset.
$h_{ii}$=E HED
Recall@20 0.3152 0.3153
NDCG@20 0.1896 0.1903
Recall@40 0.4208 0.423
NDCG@40 0.2184 0.2197
Recall@80 0.5392 0.5377
NDCG@80 0.2446 0.2453
5.6 Implementation efficiency
The number and efficiency of model parameters are also factors to be considered when evaluating the model.
For this purpose, we compared the number of parameters of the seven methods on two datasets, and the results are shown in Table 6. The majority of the work generated embeddings for users, items, and bundles. BundleGT had the highest number of parameters due to the use of multiple graph transformers. DAM had the second highest number of parameters due to the incorporation of designing attention for items. BGCN, RGNN, and CrossCBR had similar numbers of parameters. The MIDGN and UHBR models generate embeddings for users and bundles only. MIDGN has a greater number of parameters than UHBR due to the inclusion of the Intent Contrasting module. UHBR is the least parameterised model among the existing bundle recommendation algorithms. Our approach generates embeddings for users, items, and bundles, so while the number of references is higher than UHBR and MIDGN, it is lower than the other approaches, and our approach can reduce the number of references by adjusting the embedding size with some performance loss.
Table 6: A comparison of the number of parameters in different models.
Methods #Params-NetEase #Params-Youshu
DAM 18490372 5047424
BGCN 10586112 2941952
RGNN 10561538 2917378
MIDGN 4643776 836480
CrossCBR 10561280 2917120
UHBR 2649152 819904
BundleGT 31683840 8751360
HED-32 5280640 1458560
HED-64 10561280 2917120
HED-128 21122560 5834240
Similarly, we conducted a comparative analysis of the time consumption of the four methods in each period and GPU memory usage on the two datasets. The results are presented in Table 7 and Table 8. Limited by hardware, the MIGDN method is out of memory on the NetEase dataset. The UHBR model employs a smaller hypergraph and graph convolutional neural network, resulting in the fastest computational speed, with a GPU memory usage that is only slightly higher than that of CrossCBR. The CrossCBR method employs the lightGCN algorithm and a comparative learning framework with the lowest GPU memory usage, and is second only to UHBR in computational speed among the four methods. The MIDGN approach necessitates the generation of additional adjacency matrices during the graph decomposition process, which requires a significant amount of GPU memory and also results in a slower computational speed. BundleGT, on the other hand, runs slower and has a high GPU memory usage because of the graph transformer. Our method, on the other hand, exhibits different results depending on the dataset. On the Youshu dataset, our method uses a moderate amount of GPU memory and runs fast with the highest performance. With regard to the more complex NetEase dataset, it should be noted that the optimal results achieved were at the cost of high GPU memory usage and long training times. Furthermore, it is important to highlight that this level of enhancement was higher than that observed in the Youshu dataset. This is a result of the deeper mining of data, which included the construction of larger complete hypergraphs and the utilisation of dual convolutional modules in order to improve the accuracy of the recommendations. In addition, adjusting the embedding size only affects the model training time and does not reduce the GPU memory usage. Due to the limitations of our existing hardware, HED is out of memory on the iFashion dataset, and the application of our method to the industrial sector will also suffer from such problems. The recommender system utilised in industry comprises four distinct phases: recall [3], ranking [34] and re-ranking [29]. The deployment of our method in the recall phase necessitates a significant amount of memory space, whereas the deployment in the other phases circumvents the issue of hardware requirements.
Table 7: Implementation costs of different methods on NetEase datase.
Methods Training Inference GPU memory
MIDGN OOM OOM OOM
CrossCBR 5 2 2728MiB
UHBR 9 3 8412MiB
BundleGT 180.07586 177.97326 10432MiB
HED-32 27 2 23148MiB
HED-64 36 2 23148MiB
HED-128 56 4 23148MiB
Table 8: Implementation costs of different methods on Youshu datase.
Methods Training Inference GPU memory
MIDGN 7 8 5028MiB
CrossCBR 2.4 2 1642MiB
UHBR 0.71 1 2388MiB
BundleGT 29.1437 15 5392MiB
HED-32 1.69 1 3232MiB
HED-64 1.86 1 3232MiB
HED-128 2.32 1 3232MiB
6 Conclusions
In this study, the significance of creating user-item-bundle triples in bundle recommendation is examined. This paper introduces a complete hypergraph construction method and hypergraph-enhanced dual convolutional neural network (HED) models that are capable of utilising the constructed complete hypergraph. Comprehensive hypergraphs represent a methodology for elucidating the ternary relationships that emerge from user, item, and bundle interactions. This approach considers the interactions between the three entities, as well as the relationships between users and between bundles, which include interactions between each entity and itself. The HED model includes an embedding initialization layer, a dual convolutional layer, and a prediction layer. The embedding initialization layer produces embedding for every user, item, and bundle. The dual convolutional layers extract the complete hypergraph information and enhance the embedding by fusing the most important U-B graph information. The prediction layer provides recommendations based on the final embedding. Comparative experiments on Youshu and NetEase datasets demonstrate that our proposed HED model represents a new state-of-the-art approach to the field of bundle recommendation. It is notable that the HED outperforms the Youshu dataset on the more complex NetEase dataset, and that it significantly improves recommendation accuracy on both datasets. HED effectiveness was demonstrated by performing ablation studies and sensitivity analyses revealing the working mechanism. In addition, we discuss the impact of hypergraph construction on the results, analysing the efficiency of the model implementation in terms of parametric runtime and GPU memory usage.
Although HED is the recommended SOTA model for the current bundle, it is important to note that HED still has certain shortcomings. The construction of complete hypergraphs necessitates a significant allocation of memory, which consequently increases the hardware requirements for deployment of the model. Further research is required to identify methods of reducing the memory footprint and improving the learning efficiency. Furthermore, this representation of ternary relations and the mechanism of learning and enhancing the representation of ternary relations through the most important binary relations can be generalised to other problems, such as group recommendation in recommender systems.